A sonar generator on a ship produces periodic ultrasonic waves at a frequency of 2.2 MHz. The wavelength of the waves in seawater is 5.310-4 m. When the generator is directed downward, an echo is received 15.2 s later. How deep is the ocean in m?
1
Expert's answer
2013-01-10T09:27:32-0500
v=L*f
v - speed of ultrasonic waves
L - The wavelength of the waves
f - a frequency of the waves
2*l = t*v
l - a depth of the ocean
t - delay time
l = t*L*f/2 = 15.2*2.2*10^6*5.3*10^-4/2=8861.6 m